OpenStudy (anonymous):

use the substitution method to solve the given system of equations. write your answer as an order pair,(x,y) 3x -4y =13 x- 2y = 9

OpenStudy (anonymous):

(9,-11)

OpenStudy (nikita2):

x = 9 + 2y 3*(9+2y) - 4y=13 -> 18 +6y - 4y = 13 -> 2y = 13 - 18 -> y = -5/2 x = 9+ 2*(-5/2) = 4